Up to how many degrees we can heat 5litres of water, the temperature of which is + 20C, using 21 g of pretrol? Answer 44C.how? Q=46 000 000 J/kg c=4200J/kg water. m= 0,021kg petrol m=5kg water. T2=? T1= +20C. Q= q*m= 46 000 000 * 0,021= 966 000 J/kg?

1
Expert's answer
2021-05-07T17:27:46-0400

Q= q"\times"m= 46 000 000"\times"0,021= 9660000 (J)

T2 = T1 + Q / (m"\times"c) = 20 + 966000 / (5"\times"4200) = 20 + 46 = 66 oC.

Answer: T2 = 66 oC.